OpenStudy (anonymous):

Use a linear approximation (or differentials) to estimate the given number. e^0.01

OpenStudy (anonymous):

ok u just need a linear approximation of \(e^x\)

OpenStudy (anonymous):

So I should still use the f(a) + f'(a) (x-a) ?

OpenStudy (anonymous):

yes and u have\[f(x)=e^x\]and \[a=\text{whatever u want but better and simpler to use 0} \]

OpenStudy (anonymous):

ok so let me try this to see what equation i get

OpenStudy (anonymous):

iDK if what i did makes sense. I got f(0) + f'(o)(x-0) = 1 + 0(x-0) = 1

OpenStudy (anonymous):

emm but f'(0) is 1 ha?

OpenStudy (anonymous):

f'(x) = 1/x right so isnt x=0 so that makes it 1/0 = 0.

OpenStudy (anonymous):

f'(x)=e^x

OpenStudy (anonymous):

i gotta go...good night from eastern hemisphere :) think about it.

OpenStudy (anonymous):

ok tahnks!

OpenStudy (anonymous):

bye and thank u

OpenStudy (anonymous):

Its 3:35 pm here in the USA! but gnite